OpenStudy (anonymous):

The tenth term of an arithmetic sequence is 37 2 , and the second term is 5 2 . Find the first term. Can someone help?

OpenStudy (anonymous):

*37/2 and 5/2

OpenStudy (dumbcow):

for any arithmetic series An = A1 + (n-1)*d A10 = 37/2 = A1 +(10-1)*d -> 37/2 = A1 + 9d -> A1 = 37/2 - 9d To solve for d use second term and substitute in for A1 A2 = 5/2 = (37/2 -9d) +(2-1)d ->5/2 = 37/2 -8d ->(5-37)/2 = -8d ->-16 = -8d ->2=d substitute for back in for d to find A1 A1 = 37/2-9(2) =37/2 - 18 =37/2 - 36/2 = 1/2